OpenStudy (anonymous):

If f(x) = 2x + 4, find f(3). 10 6 9 4

OpenStudy (travisbrown372):

https://www.mathway.com/

undeadknight26 (undeadknight26):

to solve this questio just plug in 3 for x F(x) = 2x + 4 F(3) = 2x3+4 F(x) = 6 + 4 F(x) = 10

OpenStudy (travisbrown372):

si